OpenStudy (anonymous):

Divide, answer in a+bi form: 9/5i

OpenStudy (anonymous):

Multiply both nominator and denominator by i.

jimthompson5910 (jim_thompson5910):

\[\large \frac{9}{5i}\] \[\large \frac{9}{5i}\times\frac{i}{i}\] \[\large \frac{9i}{5i^2}\] \[\large \frac{9i}{5(-1)}\] \[\large \frac{9i}{-5}\] \[\large -\frac{9i}{5}\] So \[\large \frac{9}{5i}=-\frac{9i}{5}\]